OpenStudy (anonymous):

solve: 2/3 x -2x = 1/2 x + 11/2

OpenStudy (anonymous):

i think x = -3/5... but i'm not sure.

OpenStudy (anonymous):

2/3x-2x=1/2x+11/2 2/3x-2x-1/2x=11/2 4/6x-6/6x-3/6x=11/2 (-5/6x=11/2)-6/5 x=-66/10 or x=-33/5 or x=-6 3/5